UT 20150530 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson, BRothberg
Observer AO: A. Puglisi (remote from Italy), DMiller (remote from Tucson)
Telescope Operator: SAllenson
LBTO Support: TSargent
Watchdog: RGredel (LUCI-team)

Plan for the night

- Continue SL-checkout during first half of night

- AO-com during second half of night
- check if NCPA works as function of PA and brightnes of RS-star

- Image quality, PSF-tests if all goes well

Summary

- SL-checkout cont'd
-- if a preset fails for SL-acq. slit is put on FPU instead of turnout

- AO-com -- got excellent images in Bin=1
-- Loop opened unexpectedly several times, FLAO was not very stable

- Had problems with TCS quite a few times tonight

- obs2 freezes completey

Details

All times are UT

Files are luci2.20150530.NNNN.fits

### Calib data for SDSSJ spec taken last night

0006 0020

01:40 Steve opens for twilight flats

02:25 Sunset, clear, T = +10deg, Hum = 24%, low wind

minor adjustment to FM4 to center image

### Twilight flats
0021 0030 1 x 2.75s BrG LIR/NORM ~20...14k
0031 0040 1 x 2.75s H2 LIR/NORM ~14....9k
0041 0050 1 x 2.75s BrG LIR/NORM ~ 5....3.6k
0051 0060 1 x 2.75s H2 LIR/NORM ~ 3....2.k

0061 0065 1 x 2.75s K LIR/NORM ~12...k not dithered
0066 0070 1 x 2.75s K LIR/NORM ~9.7..9.0k
0071 0075 1 x 2.75s K LIR/NORM ~6.8k (didn't change much)

Initialize all units...OK

03:00 Steve checks pointing and collimates

Steve rested and re-flattened M2...looks better

### m1115_8 zJspec
AZ/Walth m1115_8.xml
0076 sky 0077 src 0078 slit 0079 postOffset
0080 -0.36"X offset to better center source in slit
0081 another -0.2X and also after blindOffset
0082-87

AZ/Walth m1115_8_tell.xml
-Looks like guidestar is a galaxy
-Doing a pointing check
-Resent preset -pointing is OK, guidestar is a galaxy

### NGP1301 13 01 10.34 +27 57 29.38 # HKspec
LBTO/QSO_SubMM NGP1301.xml
-04:34 - moving to this target since didn't have guidestar for m1115_8
-0088 acquisition offset
-0089 acquisition on target
-FITS headers says MASK in FPU - should have been in turnout
-88 and 89 look like darks
-92/93 new acquisition scripts after reinitializing all
-0096 - thru slit of acq star (this was blind offset) shows smeary psf
-0098-0117 spectroscopic data
-Noticed that Instrument GUI shows flexure compensation OFF
However it appears to be on for Walth_m1115 data
LBTO/QSO_SubMM NGP1301_tell_mod.xml
-made Telluric - error - couldn't find guidestar
-doing pointing check
- can't find telluric guidestar moving to next sources

LBTB/Wuyts/LUCI2
-HIP79332_mod.xml UT 06:24
-0118/0119 - acquisition script run - imaging acquisition requested
but image shows the slit. - header confirms MASK IN FPU
- this happened again after a failed preset where the guidestar
was not visible (first tonight time was above NGP 1301)
- also occurred yesterday - failed preset needed us to restart
scheduler
- Reinitialied all
- Resetting and re-committing
- 0121/122 acquisition
- manually placed in slit
- 0126-0130 Ks telluric
- after this will reconfigure manually to do zJ HK 200

-tried a 3rd time with PA=225 (3rd different PA) and new guidestar,
no guidestar found,
- preset failed caused the next execution of script to place mask in FPU
instead of to turnout
- took image of field and telluric star was there - PA was NOT what
was in preset.
- also note that we sent a PA=225 but telescope says we are at 180
- doing Telluric in track mode now (Steve sent the track mode preset
- HD 122945

>>>FW2 in unknown state UT 07:19 in move from HKspec to zJspec
- two initizlizes didn't work - need to use MCU
- moving 50 steps (197 steps between filters) into a gap
- reran initialization for FW2
- moving step by step
- ran initialization again (UT 07:23)

Filter finally at zJspec. Spectra of the telluric standard are zJspec.

Done...turning over for AO work.

07:31 Start AO-Com part of the night,
Point to A45_R8_0048, R+I about 8mag, seeing on guider about 0.9"
Put in N30 field Dot 2, FeII /clear, move in N30 camera, set focus to -60
07:42 Steve sents AO-preset, try to close the loop, optimize gain, check NCPA etc.
08:04 TCS anomaly, Steve had to restart TCS
08:08 Preset again, loop closed (yeah), seeing on guider 0.6",

  • 11:17 OBS2 X-crashed, all need to be restartet, all notes (not saved) gone...

From memory:

Using A45_R8_0048 to determine ima-quality with and without NCPA-corr, wo NCPA 153-202, with NCPA 203-252, sky's 253-262, subwindow, FeII, 0.17sec/frame
These files have been copied in the sub-folder AO

Lots of tests by Alfio and Doug on stability of FLAO and how ,uch can be offloaded without the loop be forced to open

  • Image quality NGC 6814

Did a couple of images in Br_gam, FeII and P_gam. Some images are of brilliant quality
(below 0.05" FWHM on 1min images), loop opened a few times.

Bad: If loop opens, new preset is required

Need to check out which images belong to which filter (images 152-183)

11:45 Steve is closing up

-- %USERSIG{JochenHeidt - 2015-05-30}%
Topic revision: r5 - 30 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback